1. Venus’s mass is 4.867 x 10^24 and its radius is 6,051.8 km. How much does an 81.0 kg woman weigh on Venus?
1
Expert's answer
2021-01-12T12:12:03-0500

Let's first find the acceleration due to gravity on Venus:


"g_{V}=G\dfrac{M_{V}}{R_{V}^2},""g_{V}=6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot\dfrac{4.867\cdot10^{24}\ kg}{(6.0518\cdot10^6\ m)^2}=8.86\ \dfrac{m}{s^2}."


Finally, we can find the weight of an 81.0 kg woman on the Venus:


"W=mg_V=81.0\ kg\cdot 8.86\ \dfrac{m}{s^2}=718\ N."

Answer:

"W=718\ N."